# MoodAngels: A Retrieval-augmented Multi-agent Framework for Psychiatry Diagnosis

### v1

@@ D.1.1 Self-evaluated Scales Definition
(1) CTQ (Childhood Trauma Questionnaire). A retrospective self-report inventory that measures the severity of childhood trauma across five domains: emotional, physical, and sexual abuse, as well as emotional and physical neglect.
@@
(2) DAS (Dysfunctional Attitude Scale). A self-report measure that assesses maladaptive cognitive patterns and beliefs associated with depression.
@@
(3) GAD-7 (Generalized Anxiety Disorder-7). A self-report scale used to assess the severity of generalized anxiety disorder symptoms over the past two weeks.
@@
(4) HCL-32 (Hypomania Checklist-32). A self-report questionnaire designed to identify symptoms of hypomania, often used in the assessment of bipolar disorder.
@@
(5) MDQ (Mood Disorder Questionnaire). A self-report screening tool used to identify symptoms of bipolar spectrum disorders.
@@
(6) NSSI (Non-Suicidal Self-Injury Assessment). A self-report measure designed to assess the frequency, methods, and motivations behind non-suicidal self-injurious behaviors.
@@
(7) PHQ-9 (Patient Health Questionnaire-9). A self-administered tool used to screen for and measure the severity of depression based on the DSM-5 criteria.
@@
(8) SHAPS (Snaith-Hamilton Pleasure Scale). A self-report measure designed to assess anhedonia, or the inability to experience pleasure, commonly associated with depression.
D.1.2 Clinician-rated Scales Definition
@@
(9) BPRS (Brief Psychiatric Rating Scale). A clinician-administered scale that evaluates a broad range of psychiatric symptoms, including psychosis, depression, and anxiety.
@@
(10) HAMA (Hamilton Anxiety Rating Scale). A clinician-administered scale used to measure the severity of anxiety symptoms based on observed and reported behaviors.
@@
(11) HAMD (Hamilton Depression Rating Scale). A clinician-rated tool that assesses the severity of depressive symptoms, widely used in clinical and research settings.
@@
(12) MCCB (MATRICS Consensus Cognitive Battery). A comprehensive, clinician-administered battery of tests designed to measure cognitive functioning in individuals with schizophrenia and other psychiatric disorders.
@@
(13) YMRS (Young Mania Rating Scale). A clinician-rated scale used to assess the severity of manic symptoms in individuals with bipolar disorder.
D.2 Correlation Scores Between Questions and Mood Disorder
@@
We calculated the correlation scores between each question and the total score across the 13 scales mentioned above, as well as their correlation with the presence of mood disorders, using a training set of 2,243 cases. The results are presented in Table D.2.
Table 7: Pearson Correlation (PC) coefficients and p-values between scale scores and the presence of mood disorder.
@@ 13. YMRS
Notably, for questions related to depressive mood and loss of interest, we observed variations in how similar questions were phrased across different scales. To ensure robustness, we consulted our domain expert co-author and decided to include two additional questions from the self-reported PHQ-9 depression scale: phq9_Q2 (depressed mood over the past two weeks) and phq9_Q1 (loss of interest over the past two weeks). We highlight these two question IDs in teal in Table D.2. Although their correlation scores (0.5006 and 0.5008, respectively) were slightly below the threshold of 0.5032, their inclusion provides valuable insights and enhances the diagnostic framework. This decision underscores the importance of multi-scale assessment in minimizing misjudgments of clients’ mental states, reducing the impact of short-term emotional fluctuations, and mitigating potential inaccuracies due to inconsistent responses. By repeating similar questions across self-reported and clinician-rated scales, we can achieve a more comprehensive and reliable evaluation. The detailed content and options of selected questions are listed in Table 8.
@@
Appendix E Clinical Data Processing
E.1 Medical Record Processing
@@
Effective psychiatric diagnosis relies on comprehensive and well-structured patient information. However, raw medical records often contain scattered, redundant, or sensitive details that can hinder accurate analysis. To address these challenges, we designed a systematic medical record processing pipeline that ensures data security, enhances temporal reasoning, and improves symptom extraction for LLM-based agents. This pipeline consists of four key steps: extracting essential diagnostic elements, refining present illness descriptions to prevent data leakage, integrating structured information for coherence, and an optimal step of reorganizing records into a structured format for precise symptom matching.
@@
E.2 Data Processing Prompts
Data processing steps are conducted using GPT-4o prompts followed by manual revision.
Diagnosic criteria extraction. To facilitate comparison between diagnostic criteria and medical records, please summarize the criteria for {disease name} from the DSM-5 diagnostic manual in a point-by-point format. Each point should fully describe a symptom, explicitly include reference information where necessary, and retain age or other relevant restrictions.
Below is the description of the diagnostic criteria for {disease name} from the DSM-5: {the original criteria}.
Differential Criteria Extraction. Please split the following differential criteria into two separate diagnostic criteria. Each point should fully describe a symptom, explicitly include reference information where necessary, and retain age or other relevant restrictions.
Below are the differential criteria to be split: the original criteria.
Medical Record Processing.
Step 1 $\xrightarrow{}$ Step 2: Replace specific dates in the present illness history with relative time expressions based on the patient’s admission date. Avoid using exact dates after the replacement.
Step 2 $\xrightarrow{}$ Step 3: Organize the patient’s information into a single, coherent paragraph without altering the content of the present illness section.
Step 3 $\xrightarrow{}$ Step 4: Structure the medical record by extracting and summarizing the patient’s symptoms and background. List each item separately, using "-" as a delimiter. Respond only in JSON format.
E.3 Scale Performance Description
An example of how we transform scale performances into a textual description is shown in Figure 4, and a full example is shown in Figure 5.
@@ E.3 Scale Performance Description
Appendix F Synthetic Data
F.1 Synthetic Data Construction
@@
F.2 Synthetic Data Evaluation
@@ F.2.1 Density Evaluation
Table 10: Score of column distribution. RC is short for RangeCoverage; CC is short for CategoryCoverage.
@@
F.2.2 Quality Evaluation
We evaluate synthetic data fidelity using the Alpha-Precision metric, which assesses two critical dimensions of data quality:
@@ F.2.2 Quality Evaluation
F.2.3 Machine Learning Efficiency Evaluation

### v2

@@ B.2.1 Scale Descriptions
(1) CTQ (Childhood Trauma Questionnaire). A retrospective self-report inventory that measures the severity of childhood trauma across five domains: emotional, physical, and sexual abuse, as well as emotional and physical neglect. The scale content is available at CTQ.
@@
(2) DAS (Dysfunctional Attitude Scale). A self-report measure that assesses maladaptive cognitive patterns and beliefs associated with depression. The scale content is available at DAS.
@@
(3) GAD-7 (Generalized Anxiety Disorder-7). A self-report scale used to assess the severity of generalized anxiety disorder symptoms over the past two weeks. The scale content is available at GAD-7.
@@
(4) HCL-32 (Hypomania Checklist-32). A self-report questionnaire designed to identify symptoms of hypomania, often used in the assessment of bipolar disorder. The scale content is available at HCL-32.
@@
(5) MDQ (Mood Disorder Questionnaire). A self-report screening tool used to identify symptoms of bipolar spectrum disorders. The scale content is available at MDQ.
@@
(6) NSSI (Non-Suicidal Self-Injury Assessment). A self-report measure designed to assess the frequency, methods, and motivations behind non-suicidal self-injurious behaviors. The scale content is available at NSSI.
@@
(7) PHQ-9 (Patient Health Questionnaire-9). A self-administered tool used to screen for and measure the severity of depression based on the DSM-5 criteria. The scale content is available at PHQ-9.
@@
(8) SHAPS (Snaith-Hamilton Pleasure Scale). A self-report measure designed to assess anhedonia, or the inability to experience pleasure, commonly associated with depression. The scale content is available at SHAPS.
Clinician-rated scales:
@@
(9) BPRS (Brief Psychiatric Rating Scale). A clinician-administered scale that evaluates a broad range of psychiatric symptoms, including psychosis, depression, and anxiety. The scale content is available at BPRS.
@@
(10) HAMA (Hamilton Anxiety Rating Scale). A clinician-administered scale used to measure the severity of anxiety symptoms based on observed and reported behaviors. The scale content is available at HAMA.
@@
(11) HAMD (Hamilton Depression Rating Scale). A clinician-rated tool that assesses the severity of depressive symptoms, widely used in clinical and research settings. The scale content is available at HAMD-24. In this study, we employ the Chinese version of the HAMD-24, which features a different question order compared to the English version.
@@
(12) MCCB (MATRICS Consensus Cognitive Battery). A comprehensive, clinician-administered battery of tests designed to measure cognitive functioning in individuals with schizophrenia and other psychiatric disorders. The scale content is available at MCCB.
@@
(13) YMRS (Young Mania Rating Scale). A clinician-rated scale used to assess the severity of manic symptoms in individuals with bipolar disorder. The scale content is available at YMRS.
B.2.2 Relevant Question Selection
@@
We calculated the correlation scores between each question and the total score across the 13 scales mentioned above, as well as their correlation with the presence of mood disorders, using a training set of 2,243 cases. The results are presented below:
@@ 13. YMRS
Notably, for questions related to depressive mood and loss of interest, we observed variations in how similar questions were phrased across different scales. To ensure robustness, we consulted our domain expert co-author and decided to include two additional questions from the self-reported PHQ-9 depression scale: phq9_Q2 (depressed mood over the past two weeks) and phq9_Q1 (loss of interest over the past two weeks). We highlight these two question IDs in teal. Although their correlation scores (0.5006 and 0.5008, respectively) were slightly below the threshold of 0.5032, their inclusion provides valuable insights and enhances the diagnostic framework. This decision underscores the importance of multi-scale assessment in minimizing misjudgments of clients’ mental states, reducing the impact of short-term emotional fluctuations, and mitigating potential inaccuracies due to inconsistent responses. By repeating similar questions across self-reported and clinician-rated scales, we can achieve a more comprehensive and reliable evaluation. The detailed content and options of selected questions are listed in Table 7.
@@
B.3 Retrieval Datastore
B.3.1 Mental Disorders
B.3.2 Knowledge Base Construction
Diagnostic Criteria Extraction. Symptoms in the DSM-5 are listed individually. To facilitate comparison with client records, we reframe these using GPT-4o for brevity and clarity, ensuring independent, complete criteria. We extracted symptoms for 18 common mental disorders, including all major symptoms of "bipolar and related disorders" and "depressive disorders." The extraction prompt is provided in Table 8.
Table 8: Prompts of diagnostic criteria extraction.
To facilitate comparison between diagnostic criteria and medical records, please summarize the criteria for {disease name} from the DSM-5 diagnostic manual in a point-by-point format. Each point should fully describe a symptom, explicitly include reference information where necessary, and retain age or other relevant restrictions. Below is the description of the diagnostic criteria for {disease name} from the DSM-5: {the original criteria}.
For example, one symptom of bipolar disorder is as follows:
The extracted symptom description would be:
Differential Criteria Extraction. We decompose complex differential diagnosis into distinct expressions using GPT-4o. This process ensures precise differentiation between mood disorder and other diseases. The extraction prompt is provided in Table 9.
Table 9: Prompts of differential criteria extraction.
Please split the following differential criteria into two separate diagnostic criteria. Each point should fully describe a symptom, explicitly include reference information where necessary, and retain age or other relevant restrictions. Below are the differential criteria to be split: the original criteria.
This description was refined into two distinct expressions:
B.3.3 Knowledge Base Evaluation
Appendix C Agent Details
C.1 Three Diagnostic Variants
In this section, we provide prompts of Angel.R, Angel.D and Angel.C in Table 10.
Table 10: Prompts of Angel.R, Angel.D and Angel.C.
Role: You are a psychiatry diagnosis expert. Objective: Diagnose whether the visitor {digit_id} has a mood disorder (including depression and bipolar disorder), which is characterized by depressive or manic symptoms. Constraints: 1. You may only use the following actions. 2. You must act proactively. Always keep this in mind when planning actions. Available Actions: Three variants of angels are different in actions. Angel.R: 1,2,7; Angel.D: 1,2,3,4,7; Angel.C: 1,2,5,6,7. 1. toggle_visitor_record: Retrieves structured medical records for the visitor, if available. Also extracts top-5 symptoms related to DSM-5 diagnostic criteria. Args: {"name":"digit_id", "description":"Visitor ID, must be provided in the query", "type":"int"} 2. get_scale_performances: Retrieves the visitor’s performance on the top 5% most mood disorder-related questions from psychological scales. The correlation score represents the statistical relevance to mood disorders. Args: {"name":"digit_id", "description":"Visitor ID, must be provided in the query", "type":"int"} 3. previous_cases_display: If the visitor has medical record information, this tool compares the case with past structured medical records in the database and extracts the top 5 most similar cases. Due to the personalized nature of psychiatric diagnosis, the extracted similar cases are for reference only. Args: {"name": "digit_id", "description": "Visitor ID, must be provided in the query", "type": "int"} 4. previous_scales_display: This tool compares the visitor’s scale performance with past case scale performances in the database, and extracts the top 5 most similar cases. Due to the personalized nature of psychiatric diagnosis, the extracted similar cases are for reference only. Args: {"name": "digit_id", "description": "Visitor ID, must be provided in the query", "type": "int"} 5. previous_cases_analysis: If the visitor has medical record information, this tool compares the case with past structured medical records in the database, extracts and analyzes the top 5 most similar cases. Due to the personalized nature of psychiatric diagnosis, the extracted similar cases are for reference only. Args: {"name": "digit_id", "description": "Visitor ID, must be provided in the query", "type": "int"} 6. previous_scales_analysis: This tool returns a comparison of the visitor’s scale performance with past case scale performances in the database, extracts and analyzes the top 5 most similar cases. Due to the personalized nature of psychiatric diagnosis, the extracted similar cases are for reference only. Args: {"name": "digit_id", "description": "Visitor ID, must be provided in the query", "type": "int"} 7. Finish: Completes the diagnosis process. Args:{"answer":‘yes’ or ‘no’ (indicating whether the visitor has a mood disorder)", "reasons":"A summary of key reasons supporting the decision."} Resources: You are a large language model trained on vast textual data, including factual medical knowledge. Best Practices: 1. Always refer to similar cases (medical records or scale data) for diagnosis. However, do not over-rely on past diagnoses due to the personalized nature of mental health conditions. 2. Consider all available information (medical records & scale performance). Always examine similar records (if available) and similar scale data for reference. Do not rely solely on one data source. 3. Prioritize clinical evaluation over self-reported scales in case of contradictions. Self-reported terms like "occasionally", "sometimes", and "frequently" are subjective, making it difficult for visitors to assess their condition accurately. 4. Only consider symptoms relevant to mood disorders. The visitor may have other mental disorders that are not classified as mood disorders. 5. Mood disorders include only depression and bipolar disorder, characterized by depressive or manic symptoms. Response Format: Generate a JSON string following the structure below. Do not include any extra text or explanation. {"action": {"name": "action name", "args": {"args name": "args value"}}, "thoughts": {"plan": "Briefly describe the list of short-term and long-term plans", "criticism": "Constructive self-criticism", "observation": "Summary of the current step returned to the user", "reasoning": "Reasoning behind the decision"}}
C.2 Symptom Matching
C.2.1 Medical Record Processing
@@
Effective psychiatric diagnosis relies on comprehensive and well-structured patient information. However, raw medical records often contain scattered, redundant, or sensitive details that can hinder accurate analysis. To address these challenges, we designed a systematic medical record processing pipeline that ensures data security, enhances temporal reasoning, and improves symptom extraction for LLM-based agents. This pipeline consists of four key steps: extracting essential diagnostic elements, refining present illness descriptions to prevent data leakage, integrating structured information for coherence, and an optimal step of reorganizing records into a structured format for precise symptom matching. The processing prompts are provided in Table 11.
Table 11: Prompts of medical record processing.
[…]
@@
C.2.2 Symptom Matching Formulation
C.3 Scale Performance Analysis
Client performances are converted from numeric scores to textual descriptions based on question content and options, with an example of one question provided in Figure 4, and an example of all related questions is provided in Figure 5.
@@ C.3 Scale Performance Analysis
C.4 Multi-agent Diagnosis
In this section, we provide prompts of two debate agents in Table 12 and prompts of the judge in Table 13.
Table 12: Prompts of Debate Agents.
Role: You are a psychiatry diagnosis expert. You believe the current visitor has (does not have) a mood disorder. Present your argument and persuade the opposing side. Visitor Information: This is a challenging case, and three agents have already provided their diagnoses: 1. The agent without prior case retrieval believes the visitor {has (does not have)} a mood disorder. Its reasoning: {Reasons given by Angel.R} 2. The agent who retrieved past cases but did not analyze them believes the visitor {has (does not have)} a mood disorder. Their reasoning: {Reasons given by Angel.D} 3. The agent who retrieved and analyzed past cases believes the visitor {has (does not have)} a mood disorder. Their reasoning: {Reasons given by Angel.C} Original Case Information: 1. Visitor’s Medical Record: {The visitor’s medical record in Step 3} 2. Visitor’s Scale Performances: {The visitor’s performances on selected scales} 3. Reference from Past Medical Records: {Top-5 similar medical records with analysis} 4. Reference from Past Scale Performances: {Top-5 similar scale performances with analysis} Constraints: 1. Stick to your stance. Every case presented for discussion is inherently controversial, meaning your viewpoint is strongly justified. 2. You cannot interact with physical objects. If an action is absolutely necessary, request the user to perform it. If the user refuses and no alternative exists, terminate the process to avoid wasting time and effort. Resources: 1. You are a large language model trained on extensive textual data, including factual medical knowledge. 2. All relevant visitor information is provided. Make full use of it. Best Practices: 1. Consider all available visitor information (medical records & scale performances). Examine both similar medical records and similar scale performances as references. Do not base conclusions on a single source. 2. Due to the personalized nature of psychiatric diagnosis, do not overly rely on similar case diagnoses. 3. Self-reported questionnaire results may contradict clinician evaluations. In such cases, prioritize the clinician’s assessment. Terms like "occasionally," "sometimes," and "frequently" in self-reports are subjective and may lead to misinterpretation of the visitor’s condition. 4. Only symptoms related to mood disorders should influence your diagnosis. The visitor may have other mental disorders that are not classified as mood disorders. 5. Mood disorders include only depression and bipolar disorder, characterized by depressive or manic symptoms. Response Format: Generate a JSON string in the following format. Do not include any extra text or explanation. {"response": "Your reasoning for why the visitor has (or does not have) a mood disorder, along with a rebuttal to the opposing argument.", "thoughts": {"plan": "Briefly outline short-term and long-term plans", "criticism": "Constructive self-criticism"}}
Table 13: Prompts of the Judge.
Role: You are a psychiatry diagnosis expert acting as the judge in this consultation. Your role is to determine whether each debate round should conclude and make the final diagnosis. Visitor Information: This is a challenging case, and three agents have already provided their diagnoses: Agent without prior case retrieval believes the visitor {is/isn’t, diagnosis by Angel.R} diagnosed with a mood disorder. Their reasoning: {Reasons given by Angel.R} Agent who retrieved past cases but did not analyze them believes the visitor {is/isn’t, diagnosis by Angel.D} diagnosed with a mood disorder. Their reasoning: {Reasons given by Angel.D} Agent who retrieved and analyzed past cases believes the visitor {is/isn’t, diagnosis by Angel.C} diagnosed with a mood disorder. Their reasoning: {Reasons given by Angel.C} Original Case Information: Visitor’s Medical Record: {The visitor’s medical record in Step 3} Visitor’s Scale Performances: {The visitor’s performances on selected scales} Reference from Past Medical Records: {Top-5 similar medical records with analysis} Reference from Past Scale Performances: {Top-5 similar scale performances with analysis} Constraints: Fulfill your role as a judge and base your decision solely on the arguments presented by both sides. You cannot interact with physical objects. If an action is absolutely necessary to complete the task, request the user to perform it. If the user refuses and no alternative exists, terminate the process to avoid wasting time and effort. Resources: You are a large language model trained on extensive textual data, including factual medical knowledge. All relevant visitor information is provided. Make full use of it. Best Practices: Consider all available visitor information (medical records & scale performances). Examine both similar medical records and similar scale performances as references. Do not base conclusions on a single source. Due to the personalized nature of psychiatric diagnosis, do not overly rely on similar case diagnoses. Self-reported questionnaire results may contradict clinician evaluations. In such cases, prioritize the clinician’s assessment. Terms like "occasionally," "sometimes," and "frequently" in self-reports are subjective and may lead to misinterpretation of the visitor’s condition. Only symptoms related to mood disorders should influence your diagnosis. The visitor may have other mental disorders that are not classified as mood disorders. Mood disorders include only depression and bipolar disorder, characterized by depressive or manic symptoms. Response Format: Generate a JSON string in the following format. Do not include any extra text or explanation. {"judge": "Do you believe the debate should end? Answer only ’yes’ or ’no’.", "diagnose": "Do you believe the visitor has a mood disorder? Answer only ’yes’ or ’no’.", "thoughts": { "plan": "Briefly outline short-term and long-term plans.", "criticism": "Constructive self-criticism.", "judge_reasons": "Your reasoning for why the debate should or should not end.", "reasoning": "Your reasoning for whether the visitor does or does not have a mood disorder." }}
C.4.1 Debate Details
Our debate stage is intentionally designed to simulate the classical four-stage structure commonly used in formal debates, with a minimum of 2 rounds and a maximum of 4. At the beginning of each round, the prompt guides the Pro and Con agents through a structured process:
  • Round 1 – Opening Statements: Both Pro and Con agents present their initial diagnostic positions based on the available information.
  • Round 2 – Rebuttals: Each agent critically responds to the opponent’s arguments, defending their own stance and challenging the other side’s reasoning.
  • Round 3 – Free Debate (Optional): If the Judge Agent identifies unresolved or ambiguous points after the first two rounds, it initiates a more open-ended exchange to further clarify conflicting perspectives.
  • Round 4 – Closing Statements (Optional): If the discussion remains inconclusive or highly contentious, both agents are given one final opportunity to summarize their most compelling arguments in a concise manner.
After each round, the Judge Agent evaluates whether the debate should continue, provides reasoning, and reflects on the arguments. At the end of the debate, the judge delivers a diagnosis with reasons. The example in Figure 6 demonstrates a case where the debate concluded in three rounds.
Figure 6: A debate example where only Angel.C was correct and the other two were wrong, and the Judge made a correct decision by debate.
This flexible structure allows the system to adapt the level of deliberation to the complexity of the case, balancing efficiency with thoroughness in clinical decision-making. The minimum of two rounds ensures that both sides have the opportunity to present and respond to each other’s views—this establishes a meaningful diagnostic exchange. If the Judge Agent determines that all relevant points have been sufficiently addressed by the end of Round 2, the debate can be concluded early. However, when the case involves ambiguity or conflicting symptom interpretations, the Judge may trigger additional rounds to promote deeper reasoning.
In practice, we found most cases concluded after just two rounds of debate, as the key diagnostic arguments had already been clearly presented and addressed. The remaining cases required only one additional round, with all debates concluding within three rounds. This observation supports the efficiency of our debate framework in ensuring thorough reasoning when needed, while avoiding unnecessary verbosity or repetition.
Appendix D Synthetic Data
Usage Declaration. This work and its contents, including the synthetic mental health diagnostic dataset and diagnostic agent, are provided for academic and research purposes only. None of the material constitutes medical, clinical, or therapeutic advice. The synthetic data is artificially generated and does not represent real patient information. No warranties, express or implied, are offered regarding the accuracy, reliability, or suitability of the dataset, agent, or any related content. The authors and contributors are not responsible for any errors, omissions, or consequences arising from the use of this repository.
D.1 Synthetic Data Construction
@@
D.2 Synthetic Data Evaluation
@@ D.2.1 Density Evaluation
Table 15: Score of column distribution. RC is short for RangeCoverage; $C$ is short for CategoryCoverage.
@@
D.2.2 Quality Evaluation
Automatic Evaluation We evaluate synthetic data fidelity using the Alpha-Precision metric, which assesses two critical dimensions of data quality:
@@ D.2.2 Quality Evaluation
Clincian Audit We randomly sampled 50 cases (25 real and 25 synthetic) and asked three licensed clinicians to distinguish between them. Their classification accuracy was as follows:
Table 16: Human Evaluation of MoodSyn.
Clinician Correct Real Real Misclassified Correct Synthetic Synthetic Misclassified
1 18 7 20 5
2 21 4 22 3
3 19 6 18 7
The three clinicians achieved accuracy rates of 0.80, 0.82, and 0.72 when classifying synthetic data as real, demonstrating the synthetic data’s convincing resemblance to real clinical data. We measured inter-annotator agreement using Fleiss’ Kappa, which showed moderate agreement among raters ($\kappa$ = 0.54).
D.2.3 Machine Learning Efficiency Evaluation